OpenStudy (aishanuwang):

HELP! How do I rewrite this question to solve for x? The fractions are really throwing me off..... (1/x) + (1/a) = (1/b). I need to rewrite for x = _____.

OpenStudy (phi):

you can "clear the denominators" by multiplying both sides (and all terms) by a*b*x

OpenStudy (anonymous):

x=(-ab)/(b-a)

OpenStudy (aishanuwang):

Could you write down for me each step please?

OpenStudy (anonymous):

(a+x)/xa =1/b b=xa/(x+a) b(x+a) = xa bx+ab-xa=0 x(b-a)=-ab x=(-ab)/(b-a)

OpenStudy (anonymous):

I hope you understand

OpenStudy (aishanuwang):

Thanks!

OpenStudy (anonymous):

:)